40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9116     Entry time: Fri Sep 6 23:01:08 2013
Author: Koji 
Type: Update 
Category: LSC 
Subject: Stable DRMI lock was recovered from the impact on the RF system modification 

Summary

Stable DRMI lock was recovered. The AS110 phase was adjusted. PRCL and MICH were locked with REFL33I and REFL165Q.
Still SRCL is controlled with REFL55Q.


PRMI sensing matrix

Thursday night, Jenne and I found DRMI can not be locked at all. Also the PRMI lock with REFL55 showed change in the optical gain.

In order to investigate what is happening, the PRMI sensing matrix was measured and compared with the previous one taken in the night of 8/26.

SensMat_PRMI_1000cts_580Hz_2013-08-26_235635.pngVSSensMat_PRMI_1000cts_580Hz_2013-09-06_201137.png

It shows that some signals are unchanged, some are partial change, and some are completely different.
My intuition saids something is wierd with the sensing matrix measurement.
Right now I can't trust these plots.

- Jenne and I have adjusted REFL55 demod angle so that REFL55Q has no PRCL. And I have confirmed with DTT that this is still true.
  However, the radar chart shows that REFL55Q is almost correct phase for PRCL instead of MICH.

- REFL11 shows the same amplitude and angle as before. But POX11/POY11 shows different MICH angle.

- I have rotated REFL55 demod phase and remearsured the sensing matrix. Evrything else looked same but REFL55.
  Since REFL55I&Q were not used for the control for this measurement, what we expect is to see no change of the sensing matrix and
  only see the angle of "I"&"Q" rotates. But the result was different from the expectation.

DRMI locking

Since no real info was obtained from the sensing matrix, I had to make a fight without any weapon.
After sevral hours of work, stable DRMI lock was recovered.

Basically I gave larger gains to REFL55 signals: REFL55I for SRCL was 100 instead of 1, and REFL55Q for MICH was 2 instead of 0.1.
This was enough to get a second locking. Using this short sections, I have optimized the FM triggers and the gain boosts (i.e. FM1)
as well as the mirror alignment.

Then, PRM ASS was left running during the lock. This actually stabilized the lock a lot.
This made thee lock indefinite.

The demod phase of AS110I was adjusted so that AS110Q fluctuates around zero.
In this condition, the nominal AS110I was 7300 with the whitening gain of 30dB.

Note that the AS110I&Q were also measured with PRMI. With the same phase and gains, AS110I and Q were -35,  -170, respectively.
Do we expect to have this phase shift? If I believe these numbers, the aplitude of 110MHz at the optimal phase is 173,
The ratio of AS110 between DRMI and PRMI is 7300/173 = 42. This corresponds to the ratio of the 110MHz sideband power at the AS port.
According to the wiki, this ratio shoud be ~160.

AS110I was in fact glitchy as you can see in the StripTool chart. I wonder this signal is suitable for the normalization or not.


=== SENSING ===

REFL11 -67deg / whitening gain 0dB
REFL33 -20deg / whitening gain 30dB
REFL55 45deg / whitening gain 6dB
REFL165 96deg / whitening gain 45dB

POP110 69deg whitening on / 15dB
POP22 102.2deg whitening on / 21dB
AS110 145deg whitening off / 30dB (seems to be related to AS11 whitening setting)

=== INPUT MATRIX ===

REFL11I x -0.125 => PRCL (REFL33I x 2.5 was also OK)
REFL55I x 100 => SRCL
REFL55Q x 2 => MICH (REFL165Q x 0.1 was also OK)

=== NORMALIZATION / TRIGGER ===

No normalization

Trigger settings
MICH POP22I UP:50 DOWN:10
PRCL POP22I UP:50 DOWN:10
SRCL POP22I UP:50 DOWN:25

=== SERVO FILTERS ===

MICH x -0.8 FM4/5 ON, no limitter
FM Trigger: delay 2sec, FM1 (modified from 6dB to 20dB), FM2, FM3

PRCL x +0.035 FM4/5 ON, no limitter
FM Trigger: delay 0.5sec, FM2/3/6

SRCL x -0.1 FM4/5 ON, no limitter
FM Trigger: delay 5sec, FM1, FM2

=== OUTPUT FILTERS ===

MICH => PRM -0.267 / BS +0.5

PRCL => PRM +1.0

SRCL => SRM +1.0

=== VIOLIN FILTER TRIGGER ===

delay 1sec: FM1/FM2/FM3/FM6

=== ASC/ASS ===

PRM ASC UP:50 DOWN:25
PITCH&YAW: FM1/9 (ALWAYS ON) + FM2/3 (turned on by the up-script)

PRM ASS left turned on for slow tracking

Attachment 1: DRMI.png  24 kB  Uploaded Sat Sep 7 00:37:16 2013  | Hide | Hide all
DRMI.png
ELOG V3.1.3-